### v3

@@ 3.5 Comparison Between Different Parallelisms
| Tensor Parallel | ✗ | |||
| DistriFusion | ✓ | |||
| SP-Ring | ✓ | |||
| SP-Ulysses | ✗ | |||
| PipeFusion | 2 | ✓ | ||
Compared with existing parallel solutions, PipeFusion demonstrates superior efficiency in communication, memory usage, and generation accuracy. Firstly, PipeFusion transmits between computational devices only the activations that serve as inputs and outputs for a series of consecutive transformer layers belonging to a stage, significantly reducing the communication bandwidth requirement in comparison to other methods that transmit the KV activations for all the layers. In particular, the total communication cost for PipeFusion is , not associating with the number of layers . Secondly, each device in PipeFusion stores only of the parameters belonging to its specific layers. Since the usage of stale KV for attention computation requires each device to maintain the full spatial KV for the layers on the device, the overhead is significantly smaller than that of DistriFusion and diminishes as the number of devices increases. Thirdly, PipeFusion theoretically outperforms DistriFusion considering the area of fresh activation area. As shown in Figure 4, within a single diffusion timestep, PipeFusion continuously increases the area of fresh activation as the pipeline micro-steps progress from diffusion timestep to . In contrast, throughout the entire diffusion process, DistriFusion constantly maintains one patch of fresh area out of the total patches.
@@ 4 Experiments
4.1 Setups
In our experimental setup, we deployed our trials on an 8L40-48GB (PCIe Gen4x16) cluster to evaluate three prominent DiT models.
Pixart Chen et al. (2023; 2024a): Its backbone is a transform model that encompasses 0.6B (billion) parameters. Pixart draws its architectural foundation from the original DiTs framework, with a pivotal enhancement: the integration of cross-attention modules designed to incorporate text conditions into the model’s processing.
Stable-Diffusion-3 Esser et al. (2024) (SD3-medium): which leverages a Multimodal (MM)-DiT and its backbone transformer model contains 2B parameters. As the largest open-source variant available, SD3-medium employs a 20-step FlowMatchEulerDiscreteScheduler.
Flux.1-dev BlackForestLabs (2024): Its backbone transformer model contains 12B parameters and features an enhanced MM-DiT design. Flux.1-dev also relies on a 28-step FlowMatchEulerDiscreteScheduler.
@@ 4.2 Performance Results
This section compares the performance of different parallel methods on 8L40 GPUs, connected by PCIe Gen4. This section only presents the latency of the DiT except for the latency of the VAE decoding, since VAE module is the same across different parallel methods. Since the Pixart and SD3 can utilize Classifier-Free Guidance (CFG) Ho & Salimans (2022), we additionally implemented the CFG parallel. For an input prompt, CFG parallel individually computes the forward tasks for both unconditional guidance and text guidance. It leverages inter-image parallel to perform these two tasks, collecting the results after each diffusion step is completed. For PipeFusion, we select the best latency performance by searching the patch number from 2, 4, 8, 16, 32. All figures show the average latency of 5 runs. We employ a default warmup step of 1 for both DistriFusion and PipeFusion. The software stack utilized includes PyTorch 2.4.1, CUDA Runtime 12.1.105, and diffusers 0.30.3.
@@ 4.2.1 Latency Evaluation
The highest latency parallelism is always SP-Ulysses, primarily due to the bandwidth bottleneck between CPU sockets by conducting All2All on 8 GPUs via PCIe. SP-Ring exhibits significantly lower latency compared to SP-Ulysses. An efficiency method USP Fang & Zhao (2024) can further improve SP performance by hybridizing SP-Ulysses and SP-Ring. It views the process group as a 2D mesh where the columns are SP-Ring groups and the rows are SP-Ulysses groups. USP parallelizes the attention head dimension in SP-Ulysses groups and parallelizes the sequence dimension in SP-Ring groups. USP yields a lower latency than SP-Ring, as illustrated by the ulysses=2 and ring=4 columns in the figure.
@@
Scalability: We evaluated the scalability of inference for three models by scaling from 1 to 2, 4, and 8 GPUs. Figure 8 illustrates the scalability of different parallel approaches. PipeFusion and DistriFusion both apply a single warmup iteration. We applied USP to represent SP by searching for the optimal combination of ulysses degree and ring degree to achieve the best performance. The results show that PipeFusion consistently exhibits the best scalability across all five scenarios. Given that Pixart has the smallest model size, tensor parallelization unsurprisingly exhibits the worst scalability. Except for the Pixart 4096px case, PipeFusion significantly outperforms both SP and DistriFusion. In the Pixart 4096px scenario, PipeFusion and SP exhibit similar scalability, while DistriFusion encounters out-of-memory (OOM) issues and thus lacks corresponding data. This is likely due to the exceptionally long sequence, where the communication overhead becomes less significant, thereby diminishing the advantage of PipeFusion’s communication efficiency.
@@ 4.2.2 Memory Efficiency
We collect the memory usage excluding the VAE usage, which involves the computation of the text-encoder and DiTs backbones, as shown inAs shown in Figure 1. The figure includes memory usage for ”parameters”, which encompasses both the text encoder and transformers, as well as for ”activations”, which includes activations and temporary buffers. Note that SP-Ulysses, SP-Ring and USP exhibit similar memory consumption, denoted as SP in the figure.
@@
The maximum memory usage of different parallel approaches for Pixart is depicted in Figure 9. The Pixart model consists of a 0.6B parameter DiT backbone (2.3 GB on disk) and T5-based text encoders (18GB on disk). So the memory consumption of parameters is dominated by the text encoder. DistriFusion maintains full spatial shape K, V tensors for each layer, resulting in an increase in memory consumption as the image resolutions grow. In contrast, PipeFusion only needs to store 1/ of the full spatial shape K, V tensors. Consequently, on 8GPUs, its memory footprint is comparable to that of Tensor Parallel and Sequence Parallel. Remarkably, even for tasks with an 8192px resolution, PipeFusion exhibits the lowest memory consumption.

### v4

@@ 3.4 Comparison Between Different Parallelisms
| Tensor Parallel | ||||
| DistriFusion | ||||
| SP-Ring | ||||
| SP-Ulysses | ||||
| PipeFusion | 2 | |||
Compared with existing parallel solutions, PipeFusion demonstrates superior efficiency in communication and memory usage. Firstly, PipeFusion transmits between computational devices only the activations that serve as inputs and outputs for a series of consecutive transformer layers belonging to a stage, significantly reducing the communication bandwidth requirement in comparison to other methods that transmit the KV activations for all the layers. In particular, the total communication cost for PipeFusion is , not associated with the number of layers . Secondly, each device in PipeFusion stores only of the parameters belonging to its specific layers. Since the usage of stale KV for attention computation requires each device to maintain the full spatial KV for the layers on the device, the overhead is significantly smaller than that of DistriFusion and diminishes as the number of devices increases.
@@ 4 Experiments
In our experimental setup, we deployed our trials on an 8L40-48GB (PCIe Gen4x16) cluster to evaluate three prominent DiT models: Pixart [chen2023pixart, chen2024pixart]: Its backbone is a transform model that encompasses 0.6B (billion) parameters. Pixart draws its architectural foundation from the original DiTs framework, with a pivotal enhancement: the integration of cross-attention modules designed to incorporate text conditions into the model’s processing. Stable-Diffusion-3 [esser2024scaling] (SD3-medium): which leverages a Multimodal (MM)-DiT and its backbone transformer model contains 2B parameters. As the largest open-source variant available, SD3-medium employs a 20-step FlowMatchEulerDiscreteScheduler. Flux.1-dev [blackforestlabs2023announcement]: Its backbone transformer model contains 12B parameters and features an enhanced MM-DiT design. Flux.1-dev also relies on a 28-step FlowMatchEulerDiscreteScheduler.
@@ 4.1 Performance Results
This section compares the performance of different parallel methods on 8L40 GPUs, connected by PCIe Gen4. This section only presents the latency of the DiT except for the latency of the VAE decoding, since VAE module is the same across different parallel methods. Since the Pixart and SD3 can utilize Classifier-Free Guidance (CFG) [ho2022classifier], we additionally implemented the CFG parallel. For an input prompt, CFG parallel individually computes the forward tasks for both unconditional guidance and text guidance. It leverages inter-image parallel to perform these two tasks, collecting the results after each diffusion step is completed. For PipeFusion, we select the best latency performance by searching the patch number from 2, 4, 8, 16, 32. All figures show the average latency of 5 runs. We employ a default warmup step of 1 for both DistriFusion and PipeFusion. The software stack utilized includes PyTorch 2.4.1, CUDA Runtime 12.1.105, and diffusers 0.30.3.
@@ 4.1.1 Latency Evaluation
The highest latency parallelism is always SP-Ulysses, primarily due to the bandwidth bottleneck between CPU sockets by conducting All2All on 8 GPUs via PCIe. SP-Ring exhibits significantly lower latency compared to SP-Ulysses. An efficiency method USP [fang2024unified] can further improve SP performance by hybridizing SP-Ulysses and SP-Ring. It views the process group as a 2D mesh where the columns are SP-Ring groups and the rows are SP-Ulysses groups. USP parallelizes the attention head dimension in SP-Ulysses groups and parallelizes the sequence dimension in SP-Ring groups. USP yields a lower latency than SP-Ring, as illustrated by the ulysses=2 and ring=4 columns in the figure.
@@
Scalability: We evaluated the scalability of inference for three models by scaling from 1 to 2, 4, and 8 GPUs. Figure 8 illustrates the scalability of different parallel approaches. PipeFusion and DistriFusion both apply a single warmup iteration, while for SP we adopt the USP, which searches the optimal combination of Ulysses and Ring degrees. The results show that PipeFusion consistently achieves the best scalability across all five scenarios. Given that PixArt has the smallest model size, tensor parallelization unsurprisingly exhibits the worst scalability.
At 1024px and 2048px resolutions, PipeFusion substantially outperforms both SP and DistriFusion. Even at 4096px, where DistriFusion encounters out-of-memory (OOM) issues and thus provides no valid data, PipeFusion continues to deliver strong performance: on 8L40 GPUs, it achieves 32.1s end-to-end latency versus SP’s best 37.3s, corresponding to a 1.16 speedup. While this speedup appears smaller than at lower resolutions, PipeFusion’s communication efficiency remains a clear advantage. Specifically, PipeFusion reduces communication share to only 4.6% of total latency at 4096px, compared with 17.9% for SP.
Here, the communication share is defined as
which subtracts 1/8 of the single-GPU latency from the observed 8-GPU latency. For example,
In absolute terms, SP’s communication cost is 6.69s, whereas PipeFusion lowers it to 1.49s — a 78% reduction. This shows that even in the compute-bound regime, PipeFusion robustly suppresses communication overhead.
@@ 4.1.2 Memory Efficiency
We collect the memory usage excluding the VAE usage, which involves the computation of the text-encoder and DiTs backbones, as shown inAs shown in Figure 1. The figure includes memory usage for "parameters", which encompasses both the text encoder and transformers, as well as for "activations", which includes activations and temporary buffers. Note that SP-Ulysses, SP-Ring and USP exhibit similar memory consumption, denoted as SP in the figure.
@@
The maximum memory usage of different parallel approaches for Pixart is depicted in Figure 9(a). The Pixart model consists of a 0.6B parameter DiT backbone (2.3 GB on disk) and T5-based text encoders (18GB on disk). So the memory consumption of parameters is dominated by the text encoder. DistriFusion maintains full spatial shape K, V tensors for each layer, resulting in an increase in memory consumption as the image resolutions grow. In contrast, PipeFusion only needs to store 1/ of the full spatial shape K, V tensors. Consequently, on 8GPUs, its memory footprint is comparable to that of Tensor Parallel and Sequence Parallel. Remarkably, even for tasks with an 8192px resolution, PipeFusion exhibits the lowest memory consumption.
The maximum memory usage of sequence parallel and PipeFusion for SD3-medium and Flux.1-dev is depicted in Figure 9(b). The SD3-medium consists of a 2B parameter DiTs backbone (7.8 GB on disk) and text encoders of nearly 19GB on disk. The Flux.1-dev consists of a 12B parameter DiTs backbone (23 GB on disk) and text encoders of nearly 9.1GB on disk. The memory cost of PipeFusion is also significantly less than that of sequence parallelism. For the 12B Flux.1-dev, PipeFusion significantly reduces the memory footprint of model parameters. However, PipeFusion increases the memory usage of activations due to the consumption of the KV Buffer. Nevertheless, this increase is relatively minor compared to the overall model parameter memory usage. The overall memory usage of PipeFusion is 32% and 36% of SP on 1024px and 2048px cases using Flux.1.
4.2 Quality Results
Table 2 presents the Fréchet Inception Distance (FID) [heusel2017gans] scores for PipeFusion, DistriFusion, and the Original model, evaluated on the Pixart and Flux.1 datasets. A lower FID score is indicative of superior performance in terms of image quality. The results clearly demonstrate that PipeFusion consistently outperforms DistriFusion across identical device configurations, achieving significantly lower FID scores. This finding is in direct alignment with the theoretical analysis presented in Sec. 3.3, where PipeFusion is shown to leverage fresher activations while effectively reducing the impact of stale ones, thereby enhancing overall performance.
Additionally, we provide a comprehensive visual comparison in Fig. 10 of Appendix A, where image samples generated by PipeFusion, DistriFusion, and the Original model are displayed. The images generated by PipeFusion are nearly indistinguishable from the original images to the human eye.
| Model | PixArt-XL-2-256-MS | FLUX.1[dev] | ||||||||
| Device Number | 1 | 4 | 8 | 2 | 4 | 8 | 1 | 4 | 8 | |
| w/ Orig. | - | 8.41 | 19.81 | 1.67 | 3.11 | 6.09 | - | 4.01 | 5.93 | |
5 Limitations & Discussion
While this paper has demonstrated PipeFusion’s advantages over tensor parallelism and sequence parallelism, we acknowledge that it is not a universal solution. In fact, combining PipeFusion with existing approaches into a hybrid parallelization scheme is likely the most practical direction for deployment. Such a design enables efficient scaling of DiT inference across large cross-node configurations: sequence parallelism can be exploited within NVLink-connected, high-bandwidth settings, while PipeFusion is particularly effective across inter-node, low-bandwidth environments.
A key limitation of PipeFusion lies in its dependence on multiple diffusion steps. For models with only a handful of steps—such as distilled variants like the 4-step Flux.1-schnell [flux1_schnell] or one-step diffusion methods [yin2024one, geng2025mean]—temporal redundancy is minimal, making PipeFusion less effective. However, this does not pose a major constraint in practice, as state-of-the-art DiT models generally require a substantial number of steps to deliver competitive quality.
Another promising extension is the application of PipeFusion to video generation. Architecturally, PipeFusion is already compatible with multimodal DiTs (MM-DiTs) such as Flux.1 and Stable Diffusion 3, where cross-modal conditions are injected through unified attention. Recent video models (e.g., Wan2.1 [wan2025wan], HunyuanVideo [kong2024hunyuanvideo], CogVideo [hong2022cogvideo]) adopt similar MM-DiT backbones, suggesting that PipeFusion can be applied with minimal adaptation. From the tensor-shape perspective, both image and video DiTs operate on latent tensors of shape , with video models flattening spatial and temporal dimensions into a unified sequence dimension. Because PipeFusion partitions along the sequence dimension, this structural difference does not affect applicability. From the performance perspective, video models are typically larger and involve longer sequence lengths, which amplifies communication overhead for SP. PipeFusion mitigates this overhead through parameter partitioning and pipelined execution, offering reduced memory footprint and stronger scalability. This suggests that video generation is not only a natural but also a highly favorable domain for applying PipeFusion.
6 Conclusion
We introduced PipeFusion, a patch-level pipeline parallelism strategy tailored for Diffusion Transformers. By reusing temporally redundant activations and overlapping communication with computation, PipeFusion reduces both bandwidth demands and memory footprint while preserving image quality. Experiments on Pixart, Stable Diffusion 3, and Flux.1 demonstrate consistent latency improvements over strong baselines, achieving up to 1.5× faster inference on commodity multi-GPU clusters.
Beyond efficiency gains, PipeFusion highlights a broader systems insight: temporal redundancy in diffusion processes can be systematically exploited to design new forms of parallelism. While our study focused on single-node, multi-GPU inference, the approach naturally composes with existing sequence or tensor parallelism, offering a path toward scalable deployment on large clusters. We hope this work inspires further exploration of hybrid parallel strategies that close the gap between rapidly growing generative models and the practical requirements of serving them at scale.
NeurIPS Paper Checklist
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
Answer: [Yes]
Justification: The paper clearly states its contribution of proposing PipeFusion, a novel patch-level pipeline parallelism method for efficient inference of diffusion transformer models. It highlights the key ideas of leveraging input temporal redundancy and reducing communication and memory costs compared to existing parallel methods. The claims are supported by the experimental results demonstrating state-of-the-art performance on multiple models and configurations.
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: The paper discusses limitations in Section 5, noting that PipeFusion is not well-suited for models with a limited number of diffusion steps, such as distilled models or one-step diffusion methods. It also mentions that while PipeFusion reduces communication and memory costs, it may face challenges in scenarios with extremely long sequences where communication overhead becomes less significant.
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory assumptions and proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [N/A]
Justification: The paper focuses on proposing and evaluating the PipeFusion method through experimental results, and does not include formal theoretical results, theorems, or proofs.
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental result reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Justification: The paper provides sufficient details to reproduce the main experimental results. It describes the experimental setup, including the hardware configuration (8×L40 PCIe GPUs), software stack (PyTorch, CUDA Runtime, diffusers), and the models used (Pixart, Stable-Diffusion 3, Flux.1). It also explains the methodology for implementing PipeFusion and other parallel approaches, and provides results for various image resolutions and configurations. While the code and data are not explicitly provided, the detailed descriptions allow for replication of the experiments.
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [Yes]
Justification: The paper cites the original sources for the models and datasets used, such as Pixart, Stable-Diffusion 3, and Flux.1. It also mentions the use of publicly available datasets like COCO Captions 2014 for evaluation. While it does not explicitly state the license for each asset, it is common practice in the field to assume that publicly available datasets and models are used under their respective standard licenses (e.g., CC-BY for COCO Captions). The paper does not release any new assets, so there is no need to provide licenses for derived assets.
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental setting/details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Justification:The paper cites the original sources for the models and datasets used, such as Pixart, Stable-Diffusion 3, and Flux.1. It also mentions the use of publicly available datasets like COCO Captions 2014 for evaluation. While it does not explicitly state the license for each asset, it is common practice in the field to assume that publicly available datasets and models are used under their respective standard licenses (e.g., CC-BY for COCO Captions). The paper does not release any new assets, so there is no need to provide licenses for derived assets.
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment statistical significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [No]
Justification: The paper does not report error bars, confidence intervals, or other statistical significance tests for the experiments. It is because the results is stable among tests. It presents the average latency and FID scores for multiple runs but does not provide information on the variability or statistical significance of these results.
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments compute resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: The paper provides detailed information on the computer resources used for the experiments. It specifies the hardware configuration as 8×L40 PCIe GPUs with 48GB memory each and mentions the software stack including PyTorch 2.4.1 and CUDA Runtime 12.1.105. Additionally, it reports the average latency for multiple runs, which indicates the time of execution for each experiment.
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code of ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: The research described in the paper does not involve human subjects, sensitive data, or any other elements that would require special ethical considerations according to the NeurIPS Code of Ethics. The paper focuses on improving the efficiency of diffusion transformer models for image generation, which does not pose risks related to privacy, discrimination, security, or other ethical concerns outlined in the guidelines.
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [No]
Justification: The paper does not discuss the broader societal impacts of the work performed. While it focuses on improving the efficiency of diffusion transformer models for image generation, it does not address potential positive or negative societal impacts. The research is foundational and does not directly tie to specific applications or deployments that could have significant societal consequences.
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [N/A]
Justification: The paper does not release any datasets or models that pose a high risk for misuse. It focuses on the methodology and experimental results of the PipeFusion approach without distributing any models or datasets that could be misused.
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: The paper cites the original sources for the models and datasets used, such as Pixart, Stable-Diffusion 3, and Flux.1. It also mentions the use of publicly available datasets like COCO Captions 2014 for evaluation. While it does not explicitly state the license for each asset, it is common practice in the field to assume that publicly available datasets and models are used under their respective standard licenses (e.g., CC-BY for COCO Captions). The paper does not release any new assets, so there is no need to provide licenses for derived assets.
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [N/A]
Justification: The paper does not introduce or release any new assets such as datasets, models, or code. Therefore, there is no need to document or provide documentation for new assets.
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and research with human subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [N/A]
Justification: The paper does not involve any crowdsourcing experiments or research with human subjects. It focuses on the technical aspects of parallelizing diffusion transformer models for image generation, which does not require human participation or data collection.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional review board (IRB) approvals or equivalent for research with human subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [N/A]
Justification: The paper does not involve any research with human subjects, so IRB approvals or equivalent reviews are not applicable.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
16.
Declaration of LLM usage
Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
Answer: [No]
Justification: The paper does not describe the usage of Large Language Models (LLMs) as an important, original, or non-standard component of the core methods. The research focuses on the development and evaluation of the PipeFusion method for parallelizing diffusion transformer models, which is a technical contribution unrelated to LLMs.
Guidelines:
The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Image Quality
In Figure 10, we selected a few sample images generated using prompts from DistriFusion [li2023distrifusion] to help readers visually assess generation quality. The images generated by PipeFusion are nearly indistinguishable from the original images to the human eye, across various configurations of the number of patches and devices. Note that the FID evaluation results are independent of the specific prompts used for the visual demonstrations.